5r-2s=37
2r+5s=38
Solve using the elimination method?

Answer by stanbon(75887)

5r-2s=37
2r+5s=38
-----------------
Multiply 1st by 2; Multiply 2nd by 5.
--------------------
10r-4s = 74
10r+25s = 190
------------------------
Subtract 1st from 2nd and solve for "s":
29s = 116
s = 4
----
Substitute for "s" and solve for "r:
5r - 2*4 = 37
5r = 45
r = 9
=======================
Cheers,
Stan H.